OpenStudy (anonymous):

make a scale drawing of a rectangular schoolhall which is 50m long and 30m wide using a scale of 1cm to 5m, use your scale drawing to find the actual distance between the opposite corners of the hall.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

no

OpenStudy (skullpatrol):

Do you know what a "scale drawing" is?

OpenStudy (anonymous):

no im afraid i dont

OpenStudy (skullpatrol):

Do you know what a "drawing" is?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

Good :) Let's start there. What is a "drawing"?

OpenStudy (anonymous):

i was under the impression that i type ques and the comp gives answers...:)

OpenStudy (anonymous):

hello?

OpenStudy (skullpatrol):

Hi.

OpenStudy (anonymous):

sorry there was something wrong with my internet connection.

OpenStudy (skullpatrol):

np

OpenStudy (anonymous):

Don't we have the height of the hall?

OpenStudy (anonymous):

so i draw a scale drawing of the following measurements.. then?

OpenStudy (anonymous):

50m long and 30m wide

OpenStudy (skullpatrol):

As the question states: "use your scale drawing to find the actual distance "

OpenStudy (skullpatrol):

So that means you would measure it off the diagram.

OpenStudy (anonymous):

no thats not what it means

OpenStudy (anonymous):

but thanku.

OpenStudy (anonymous):

Yeah,but that gives you only 2 out of the 3 dimensions of the hall. To draw the 3D diagram,we would need the height

OpenStudy (skullpatrol):

@najda What does it mean to you?

OpenStudy (skullpatrol):

@AishwaryaHB strictly speaking the "opposite" corner is from the ground to the roof, but since we are not given the height we are left with a two dimensional "opposite".

OpenStudy (anonymous):

In that case,your answer would be 58.3m,or something around that value

OpenStudy (skullpatrol):

But the question asks to use the scale drawing.

OpenStudy (skullpatrol):

make a scale drawing of a rectangular school hall which is 50 m long and 30 m wide using a scale of 1 cm to 5 m, use your scale drawing to find the actual distance between the opposite corners of the hall.

OpenStudy (anonymous):

yeah,so make 1! according to the scale given,1cm=50 m. so,in the drawing,your length would be=50/5=10cm and your width would be=30/5=6cm. It's a rectangular hall,so draw a rectangle with length 10cm and width 6cm. The length of the diagonal of this rectangle will represent the distance between any 2 opposite corners of the hall.

OpenStudy (anonymous):

The length of the diagonal is ~11.6cm since 1cm=5m, 11.6cm=5*11.6m =58m